OpenStudy (anonymous):

4-7). Solve the following systems by solving one of the equations for x or y and then using the substitution method. x-4y=-4 x-3y=0 what is x? what is y?

OpenStudy (anonymous):

x = 3y 3y -4y = -4 -y=-4 y=4 x=3y x=12